OpenStudy (solomonzelman):

OK, now I won't post a smile, if moderators think it is worse than a dot or a comma. I need help with an SAT question. If f(3) = 15 and f(5) = 45, which of the following could be f(x)? A. 4x + 3 B. 2x² – 2x C. 2x² - x D. 2x² - 5 E. 5x²

OpenStudy (shamil98):

plug in the values.

OpenStudy (shamil98):

4(3) + 3 = 15 4(5) + 3 = 23 so you can cross off A.

OpenStudy (solomonzelman):

cross off A?

OpenStudy (solomonzelman):

I see!

OpenStudy (shamil98):

2(3)^2 - 2(3) = 18 -6 = 12 so it's not B either.

OpenStudy (solomonzelman):

and just keep plugging in points till i get the right answer?

OpenStudy (shamil98):

yes

OpenStudy (solomonzelman):

C

OpenStudy (shamil98):

its obviously not D either

OpenStudy (solomonzelman):

C

OpenStudy (anonymous):

process of elimination is best sometimes when you are given choices

OpenStudy (shamil98):

yeah

OpenStudy (solomonzelman):

Yeah! Someone helped! @shamil98, ty!

OpenStudy (shamil98):

2(5)^2 - 5 = 45 2(3)^2 - 5= 15 yeah, np.

OpenStudy (solomonzelman):

I have a different question, why did a moderator issue me when I posted a smile?

OpenStudy (shamil98):

Was there a question in the replies?

OpenStudy (shamil98):

It could be because it was irrelevant to the Mathmematics section, off-topic.

OpenStudy (solomonzelman):

No, I pereviously posted a q, wanting to ask this, but Arving came, whatever his name is, and closed it.

OpenStudy (wolf1728):

I didn't know you could post a smilie. I always use :-)

OpenStudy (solomonzelman):

some people post dots, and why is a dot more relevant than smile?

OpenStudy (solomonzelman):

you could (•‿•)

OpenStudy (shamil98):

For that then just ask that moderator or another mod about it.

OpenStudy (solomonzelman):

Yeah, and they are going to warm me of being suspicious. then will find something to stick to, idc, you helped me, that;s good enough.

OpenStudy (wolf1728):

You can always try this ☺

OpenStudy (solomonzelman):

Cool

OpenStudy (usukidoll):

you could plug x = 5 and x = 3 into the equations and see if it comes up with the result

OpenStudy (solomonzelman):

Everyone! I just fanned @shamil98 andf @wolf1728! You deserve it!!!

OpenStudy (solomonzelman):

that 2

OpenStudy (wolf1728):

Well thanks !!! ☺

OpenStudy (solomonzelman):

Anytime, ty!

OpenStudy (anonymous):

What about that one^^

OpenStudy (wolf1728):

By the way to get the smilie face hold the alt key and then press 1

OpenStudy (solomonzelman):

No, doesn't work

OpenStudy (usukidoll):

I agree with that

OpenStudy (anonymous):

it doesnt work on my kb also

OpenStudy (solomonzelman):

maybe, cuz I have windows8, dk how would that make a difference.

OpenStudy (usukidoll):

I have windows 7

OpenStudy (anonymous):

i am on ubuntu linux

OpenStudy (anonymous):

copy it from google <---------- best method

OpenStudy (solomonzelman):

Yeah. but it is kind of small!

OpenStudy (wolf1728):

I have windows 7 and it does work for me - then again it's Microsoft™ so maybe it has to do with an update to Windows 98

OpenStudy (solomonzelman):

@╰☆╮Openstudier╰☆╮, that is a cool attachment!

OpenStudy (solomonzelman):

windows98? LMFAO

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

windows 98 hahah

OpenStudy (solomonzelman):

even win 9 is still LMF.....

OpenStudy (solomonzelman):

unless you are living 5-6 yrs later, than yeah.

OpenStudy (solomonzelman):

sorry

OpenStudy (anonymous):

lol

OpenStudy (wolf1728):

I just brought up Win 98 because Microsoft is so messed up maybe the problem could have started way back then

OpenStudy (solomonzelman):

Win 98? I've never heard of that

OpenStudy (wolf1728):

Yeah I think it goes back to the late 1990's - it was notorious for crashing and you woluld lose everything you had typed since your last save. Good old Microsoft

OpenStudy (wolf1728):

When it crashed, you'd get the infamous blue screen (I bet Microsoft invented that !!!)